# Hybrid-z: Enhancing Kilo-Degree Survey bright galaxy sample photometric redshifts with deep learning

### v1

@@ 2.3 GAMA Spectroscopic Data
KiDS DR4 fully overlaps with four GAMA fields (all but G02). Of these, the equatorial ones present the largest flux-limited spectroscopic completeness, originally estimated as 98.5% at $r_{\mathrm{SDSS}}<19.8$ mag (Liske et al., 2015) but subsequently revised to 98% at $r_{\mathrm{KiDS}}<19.58$ (Driver et al., 2022) after ingestion of KiDS photometry into the GAMA database (Bellstedt et al., 2020). Following the results of B21, where the addition of the shallower and less complete G23 data did not lead to improvement in photo-$z$ estimates, also here we used only the equatorial fields to train the model, and it is tested on both equatorial and G23 fields. Similarly, other datasets such as for instance SDSS or DESI Early Data Release, are not sufficiently complete at our flux limit of $r<20$ mag to provide useful training sets for the overall galaxy sample (see e.g. Jalan et al. 2024). Some of them are however useful for a posteriori tests of our photo-z model, and this is discussed in Sec. 5.2.
In this work, we employ the GAMA-II spectroscopic redshift catalog of galaxies from the final GAMA DR4 (Driver et al., 2022) [^3] http://www.gama-survey.org/dr4. For secure redshifts, we select galaxies with a normalized quality parameter $NQ\geq 3$ (see Liske et al., 2015, for details) and $z>0.001$. We have identified galaxies within KiDS tiles that are shared between the GAMA and KiDS-Bright samples based on their right ascension and declination coordinates, assuming a 1” matching radius between KiDS and GAMA. Then we made cutouts of size $7.2^{\prime\prime}\times 7.2^{\prime\prime}$ from KiDS images with these galaxies positioned at the center.The final cutout catalog contains $\sim 173$k galaxies in the equatorial fields and $\sim 125$k of these are used to train the model. The model is validated on $\sim 26$k galaxies in the equatorial fields, and the remaining data are used for testing. Finally, it is applied to the entire KiDS-Bright Sample of about 1.2 million galaxies.
@@ 3.1 Convolutional Neural Networks
A Convolutional Neural Network is a sequence of layers: convolutional, pooling, and fully-connected ones. The convolution takes place between input images and a small matrix of weights called a kernel or filter. Initially, the weights are small random values.
In CNN, the learning is hierarchical. The first convolutional layers are responsible for extracting the lower-level features of images, such as edges, corners, and textures (Goodfellow et al., 2016). The kernels adjust their weights to extract these features. The filters in the following layers help to learn more complex features of input data. The network is trained by non-linear optimization of weights and biases through a gradient-descent algorithm. Input data is transformed by convolution operation, producing linear activation.
@@ 3.2 Training
The training of a neural network involves several key steps and considerations. Initially, the dataset is divided into three subsets: training, validation, and testing set, in our case in a ratio of 70:15:15. The validation set serves to monitor the performance of the network during the training process.
@@ 3.3 Data Augmentation
One of the solutions for over- and underfitting is generating a large training dataset using data augmentation. Data augmentation is commonly used in ML, particularly in the context of image classification and computer vision tasks (Shorten & Khoshgoftaar, 2019). It involves generating additional training data by applying various transformations (rotation, width and height shift, flipping, etc.) to the existing training samples. After learning a certain pattern in an image, CNN can recognize it anywhere due to its translational invariance.
@@ 3.4 Inception Module
Figure 1: The Inception module used in this study to estimate the photo-$z$. The input layer has size (36,36,4). Conv2D is a two-dimensional convolution layer, with the kernel size specified in brackets. Average Pooling 2D is a two-dimensional pooling operation that uses a kernel of size $3\times 3$. Each operation is represented using boxes of distinct colors. Concatenation is the combined feature maps from parallel convolutions.
In this section, we discuss a CNN architecture named ‘Inception’ that we employ in photo-$z$ prediction.
Increasing the depth of the network improves the performance at a cost of high computation time. Inception, a deep CNN architecture developed for image detection and classification tasks. It was first used in GoogLeNet architecture (Szegedy et al., 2015). The smaller number of weights, and biases compared to its predecessors reduces the computation time and makes it appropriate for big-data handling. There are various versions of Inception available, differing by the use of regularization, reduction of overfitting when the training sample is limited, and the inclusion of additional DL architectures such as Resnet (He et al., 2015) in the Inception module. In our model, we used the Inception v1 module illustrated in Fig. 1. We also tested other versions and we found that Inception v1 gives the best performance for our dataset.
Choosing the right kernel is very important for feature extraction. On the one hand, a larger kernel is suitable for images where the information is distributed globally. However, a smaller-sized kernel is good for locally distributed information extraction. This selection becomes highly important as the galaxies in our sample have a wide range of apparent sizes, and hence, our data set contains both kinds of distributions. We used both larger and smaller-sized kernels in the Inception module. This is shown in Fig. 1.
@@ 3.5 Metrics
In this Section, we discuss the metrics to evaluate the performance of our model and photo-zs. Here we discriminate the two types of metrics into those which are evaluated and optimized when building the model (during training and validation), and those computed a posteriori to quantify the behavior of the photo-zs. For the former, we mainly used three metrics: Mean Squared Error (MSE), Mean Absolute Error (MAE), and the $R$ squared error ($R^{2}$).
MSE and MAE are defined as follows:
@@
$\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|z_{i}-\hat{z}_{i}|$ (4)
Here, $n$ is the number of samples used for training, $z_{i}$ is the predicted value and $\hat{z}_{i}$ is the true value.
$R^{2}$ is a statistical measure that shows how well a model predicts the outcome in a regression analysis. It is the proportion between the variance explained by the model and the total variance:
@@
Here, $\bar{z}$ is the mean of the spectroscopic redshift values from the GAMA×$\times$×KiDS cross-match. For an ideal case, the value of this metric would be unity.
MSE is more sensitive to outliers than MAE. Due to the squaring part, MSE puts more weight on larger errors. MAE, on the other hand, treats all errors with equal importance, which can be advantageous when the dataset contains extreme values or noisy observations – which is common for galaxy images. Therefore, MSE is more appropriate for learning outliers, and MAE is better for ignoring them. To include the benefits of both these loss functions, we use the Huber (1964) loss function, similarly as in the previous KiDS analysis by Li et al. (2022). This function is quadratic when the absolute error is small, and linear when the absolute error exceeds a threshold $\delta$. This makes it robust to outliers because the impact of large errors is reduced compared to using a purely quadratic loss function like MSE. The Huber loss is defined as
$\textit{L}(z,\hat{z})=\cases{}{1}{2}(z-\hat{z})^{2},&\text{if }|z-\hat{z}|\leq\delta,\\ \delta(|z-\hat{z}|-\frac{\delta}{2}),\text{otherwise}.{}$ (6)
@@
We evaluate the resulting photo-z performance using the following standard statistics:
@@
    $\delta z=z_{\mathrm{phot}}-z_{\mathrm{spec}},$ (7)
@@
    $\Delta z=\frac{\delta z}{1+z_{\mathrm{spec}}},$ (8)
  • standard deviation of normalized bias, $\sigma_{\Delta z}$,
@@ 4 Photometric redshift model
We used the Rectified linear unit (Relu, Jentzen et al., 2023) as the activation function in all the layers except in the output one where the sigmoid function is used, which enforces all the predictions to lie in the range $10<z_{\mathrm{phot}}<$ 1. We also tried other activation functions such as leaky Relu (Jentzen et al., 2023), peaky Relu, softmax, swish, and tanh. In our case, the sigmoid function gives the best results in our redshift range where practically all the galaxies have $z<1$ due to the $r<20$ mag flux limit. For instance, in the entire GAMA spectroscopic sample, objects with $z>1$ constitute less than 0.1% of the total and these are typically very bright and rare AGNs. For such sparse sources, empirical methods such as ours would not be able to render reliable redshift predictions unless some special approach to anomaly handling is taken. We therefore sacrifice a very small number of objects that lie at true $z>1$ to have $z_{\mathrm{phot}}<1$, avoiding instead gross redshift overestimation for others, which could happen if the model had more freedom. The latter is for instance the case for ANNz2, where some of the photo-zs from B21 are predicted significantly above unity.
Saturation of the sigmoid function happens for large or small input values, meaning its output becomes close to 1 or 0, respectively. When the input to the sigmoid function is far from zero, the gradient of the sigmoid function becomes very small, leading to vanishing gradients during backpropagation. Normalizing input data to the range $[0,1]$ ensures that the inputs to the sigmoid function are more likely to be in the region where the sigmoid function has a non-saturated gradient, making the training process more stable. Relu does not suffer from the vanishing gradient problem in the same way as the sigmoid function, but having input in the $[0,1]$ range helps maintain stable gradients.
Figure 2: Architecture of the Hybrid-z model, which employs 4-band KiDS galaxy images together with 9-band KiDS+VIKING magnitudes for photo-z derivation. For the CNN part (left-hand side), symbols used are as in Fig. 1 while Inception modules are represented as hexagons. In the fully-connected part (right-hand side), dense layers are given as blue rectangles and the number of neurons is shown in parentheses.
@@
The features extracted by the convolutional and average pooling layers are passed through the Inception module before reaching the final dense layers. We found that the performance of the model improves with the inclusion of Inception, with notable improvement found in the reduction of the loss function value. We conducted tests with different numbers of Inception modules, ranging from three to five. Our findings suggest that using four Inception modules with varying numbers of filters is optimal for our data. The first inception module receives input from the Average Pooling 2D layer, while each subsequent inception module takes its input from the output of the preceding inception module. The initial inception modules extract low-level features and the final inception layers capture the global patterns in the image.
Figure 3: Performance of the Huber loss function during training and validation stage of the Hybrid-z.
Similarly as in e.g. Li et al. (2022) or Henghes et al. (2022), for our photo-z model, we combined two different types of networks, ‘shallow’ (fully-connected) ANN, that we will denote as ordinary neural networks (ONN) and CNN, hence the name Hybrid-z. An important aspect of the Hybrid-z model is the concatenation step. The nine-band magnitudes of galaxies are the additional information for the network. These are processed by dense layers. The flattened feature map and ONN output are combined via depth-wise concatenation. Then this concatenated information is passed through final layers for the prediction of photo-zs.
In Fig. 3, the performance of the Huber loss function for the Hybrid-z is shown. The addition of supplementary information enhances the model’s performance with respect to using only 4-band images. We find that for this model, we can achieve $R^{2}$ value greater than 0.93. We used early stopping criteria for determining the number of epochs. The training will automatically stop if there is no improvement in the validation loss.
@@ 5.1 Hybrid-z performance on test data
    General test data from the KiDS-DR4 Bright sample cross-matched with GAMA in the equatorial fields, including both clean and contaminated galaxies.
@@
Figure 4: Comparison of the true spectroscopic redshifts with the photometric ones derived with our framework Hybrid-z, using 4-band KiDS images and 9-band KiDS+VIKING magnitudes. The test data were taken from a subset of GAMA cross-matched with KiDS. The thick red line depicts the median while the thin lines show the scatter (SMAD) around the median.
Figure 5: Photometric redshift errors of Hybrid-z as a function of four quantities, from left to right: (a) photometric redshift as derived in this work; (b) true spectroscopic redshift from the clean sample; (c) apparent $r$-band magnitude; and (d) observed $u-g$ color. The thick solid red line represents the running median, while the thin red lines illustrate the scatter (SMAD), based on a blind test sample derived from GAMA.
The comparison of the main statistics of Hybrid-z vs. ANNz2 results are provided in Table 1. Our new model employing jointly optical imaging and KiDS+VIKING magnitudes performs clearly better than ANNz2 which used 9-band magnitudes. In the first block of the Table, we provide statistics of a general test sample, which consists of GAMA galaxies cross-matched with KiDS-Bright including both ‘clean’ (masked=0) and ‘contaminated’ (masked=1) data, where the ‘masked’ flag was derived in B21 based on the KiDS-internal ‘MASK’ bit-wise column. The compared methods have comparable mean residuals – that is, they are similarly accurate – but Hybrid-z outperforms ANNz2 in its photo-$z$ precision, quantified as the scatter in $\Delta z$. Both in terms of standard deviation and SMAD, our combined CNN+ONN approach performs better, be it for the overall or clean (masked=0) test data. Hybrid-z provides about 20% smaller scatter than the ANNz2-based derivations of B21, where 9-band magnitudes were used. In the case of the fiducial clean data that are recommended for science, we achieve SMAD$(\Delta z)\simeq 0.014$, as compared to 0.018 in B21. Our result is at the same level as that by Treyer et al. (2024) where CNNs were used for photo-$z$s based on SDSS images at similar depths as in our case.
@@
In panels (c) and (d) of Fig. 5 we illustrate the behavior of photo-$z$ errors of our Hybrid-z model as a function of two observables: the $r$-band apparent magnitude and the observed $u-g$ color. The former is shown for the ‘auto’ flux measurement, which approximates the total light of a galaxy and was used by B21 to select the flux-limited KiDS-Bright sample. We observe slightly growing scatter (thin red lines indicating SMAD in Fig. 5 (c)) at the faintest end of the sample ($r\lesssim 20$). Otherwise, however, photo-$z$s are stable as a function of magnitude, with practically zero bias at most of the range. The $u-g$ color shown in panel (d) can be used to split galaxies into red and blue populations (similarly as $u-r$, Strateva et al. 2001). Indeed, we see the bimodality in the plot, with bluer galaxies to the left and redder to the right. As expected for photo-$z$s, the former has a wider scatter with respect to the true value than the latter. The difference is however not substantial and again, the median biases for the whole range of the $u-g$ color are very close to 0.
@@
We further quantify the performance of Hybrid-z for blue and red galaxies in the two bottom blocks of Table 1. Here, for a direct comparison between this work and B21, the selection of the two galaxy types was based on their intrinsic rather than observed properties, namely via their positioning on the absolute $r$-band magnitude versus rest-frame $u-g$ color diagram [^4] More specifically, we cross-matched our sample with galaxies labeled as red and blue in KiDS-Bright DR4. as discussed in B21. Similarly as for the full sample, also for these subsets, Hybrid-z provides considerable improvement in precision over ANNz2. Interestingly, this gain is relatively larger for blue galaxies (by $\sim 22\%$) than for red ones ($\sim 17\%$). This is consistent with the fact that, as compared to the approaches relying solely on summary measurements like magnitudes, the DL-based methods, which directly extract features from images to estimate redshifts, are expected to show larger improvement for galaxies with intricate morphologies such as spirals, that typically have bluer colors (e.g. Schuldt et al., 2021; Treyer et al., 2024). In contrast, for red galaxies – typically ellipticals with fewer features in images – the reduction in scatter for Hybrid-z is smaller than for the overall sample, although we would like to emphasize that it still does perform better than ANNz2 for the same galaxy selection. In addition, it is worth noting that SMAD($\Delta z$) of 0.0154 for blue galaxies as obtained by Hybrid-z is even smaller than the same statistic of ANNz2 for red ones (cf. blocks #3 and #4 of Table 1).
@@
Figure 6: Comparison of photometric redshift error distributions between Hybrid-z (blue, this work) and ANNz2 (B21, red), shown as normalized counts based on the test sample results. Solid lines illustrate the best-fit Gaussian (with $\sigma=0.0145$ and $\mu=2.56\times 10^{-4}$, yellow) and generalized Lorentzian (Eq. 10, with $a=2.332$ and $s=0.0116$, black). The inset compares the histogram of Hybrid-z photo-z residuals with the fitted models in logscale of the $y$-axis.
@@
The final comparison between ANNz2 and Hybrid-z performance is provided in Fig. 6, where we show the distributions of photo-$z$ errors $\Delta z$ as calibrated on the GAMA test data. As discussed in B21, following earlier work (e.g. Bilicki et al., 2014), such a distribution presents non-Gaussian features in the wings and is therefore better fitted with a ‘generalized Lorentzian’ of the form [^5] It is interesting to note that also spectroscopic redshifts may present Lorentizan rather than Gaussian uncertainties, e.g. Yu et al. (2024).
@@
To summarize this sub-section, the Hybrid-z model using CNNs on KiDS optical images together with ONNs on KiDS+VIKING magnitudes provides considerably better results than the previous 9-band ANNz2 derivations of B21. The improvement from ANNz2 to Hybrid-z is at the level of 20% reduction in photo-$z$ scatter (better precision) while maintaining almost zero bias (very good accuracy). Thanks to our new model, we observe smaller photo-$z$ errors for various subsamples of test data, and most remarkably for blue galaxies, where SMAD($\Delta z$) is lowered by about 22% from what was obtained in B21. As such, these are state-of-the-art results for the KiDS-Bright galaxies and in line with independent derivations for SDSS at the same depth (Treyer et al., 2024).

### v2

@@ 2.3 GAMA spectroscopic data
KiDS DR4 fully overlaps with four GAMA fields (all but G02). Of these, the equatorial ones present the largest flux-limited spectroscopic completeness, originally estimated as 98.5% at $r_{\mathrm{SDSS}}<19.8$ mag (Liske et al., 2015) but subsequently revised to 98% at $r_{\mathrm{KiDS}}<19.58$ (Driver et al., 2022) after ingestion of KiDS photometry into the GAMA database (Bellstedt et al., 2020). Following the results of B21, where the addition of the shallower and less complete G23 data did not lead to improvement in photo-$z$ estimates, also here we used only the equatorial fields to train the model, and it is tested on both equatorial and G23 fields. Similarly, other datasets such as for instance SDSS or DESI Early Data Release, are not sufficiently complete at our flux limit of $r<20$ mag to provide useful training sets for the overall galaxy sample (see, e.g., Jalan et al. 2024). Some of them, however, are useful for a posteriori tests of our photo-z model, and this is discussed in Sec. 5.2.
In this work, we employ the GAMA-II spectroscopic redshift catalog of galaxies from the final GAMA DR4 (Driver et al., 2022). [^3] http://www.gama-survey.org/dr4 For secure redshifts, we select galaxies with a normalized quality parameter $NQ\geq 3$ (see Liske et al., 2015, for details) and $z>0.001$. We have identified galaxies within KiDS tiles that are shared between the GAMA and KiDS-Bright samples based on their right ascension and declination coordinates, assuming a 1” matching radius between KiDS and GAMA. Then we made cutouts of size $7.2^{\prime\prime}\times 7.2^{\prime\prime}$ from KiDS images with these galaxies positioned at the center.
@@ 3.1 Convolutional neural networks
A CNN is a sequence of layers: convolutional, pooling, and fully connected layers. The convolution takes place between input images and a small matrix of weights called a kernel or filter. Initially, the weights are small random values. In a CNN, the learning is hierarchical. The first convolutional layers are responsible for extracting the lower-level features of images, such as edges, corners, and textures (Goodfellow et al., 2016). The kernels adjust their weights to extract these features. The filters in the following layers help integrate more complex features of input data. The network is trained by non-linear optimization of weights and biases through a gradient-descent algorithm. Input data is transformed by convolution operation, producing linear activation.
@@ 3.2 Training
The training of a neural network involves several key steps and considerations. Initially, the dataset is divided into three subsets in a random manner using scikit-learn python library: training, validation, and testing set, in our case in a ratio of 70:15:15. The validation set serves to monitor the performance of the network during the training process. Also, there is no selection bias in magnitudes or redshifts in any subset. The final cutout catalog contains $\sim 173$k galaxies in the equatorial fields and $\sim 125$k of these are used to train the model. The model is validated on $\sim 26$k galaxies in the equatorial fields, and the remaining data are used for testing. Finally, it is applied to the entire KiDS-DR4 Bright sample of about 1.2 million galaxies.
@@ 3.3 Data augmentation
One of the solutions for over- and underfitting is generating a large training dataset using data augmentation. Data augmentation is commonly used in ML, particularly in the context of image classification and computer vision tasks (Shorten & Khoshgoftaar, 2019). It involves generating additional training data by applying various transformations (rotation, width and height shift, flipping, etc.) to the existing training samples. After learning a certain pattern in an image, a CNN can recognize it anywhere due to its translational invariance.
@@ 3.4 Inception module
Figure 1: Inception module used in this study to estimate the photo-$z$. The input layer has size (36,36,4). Conv2D is a two-dimensional convolution layer, with the kernel size specified in brackets. Average Pooling 2D is a two-dimensional pooling operation that uses a kernel of size $3\times 3$. Each operation is represented using boxes of distinct colors. Concatenation is the combined feature maps from parallel convolutions.
In this section, we discuss a CNN architecture referred to as ”inception” that we employ in photo-$z$ prediction. Increasing the depth of the network improves the performance at a cost of high computation time. Inception, a deep CNN architecture developed for image detection and classification tasks. It was first used in GoogLeNet architecture (Szegedy et al., 2015). The smaller number of weights, and biases compared to its predecessors reduces the computation time and makes it appropriate for big-data handling. There are various versions of Inception available, differing by the use of regularization, reduction of overfitting when the training sample is limited, and the inclusion of additional DL architectures such as Resnet (He et al., 2015) in the Inception module. In our model, we used the Inception v1 module illustrated in Fig. 1. We also tested other versions and we found that Inception v1 gives the best performance for our dataset.
Choosing the right kernel is very important for feature extraction. On the one hand, a larger kernel is suitable for images where the information is distributed globally. However, a smaller-sized kernel is good for locally distributed information extraction. This selection becomes highly important as the galaxies in our sample have a wide range of apparent sizes, and hence our dataset contains both kinds of distributions. We used both larger and smaller-sized kernels in the Inception module. This is shown in Fig. 1.
@@ 3.5 Metrics
In this Section, we discuss the metrics to evaluate the performance of our model and photo-zs. Here we discriminate the two types of metrics into those which are evaluated and optimized when building the model (during training and validation), and those computed a posteriori to quantify the behavior of the photo-zs. For the former, we mainly used three metrics: mean squared error (MSE), mean absolute error (MAE), and the $R$ squared error ($R^{2}$). The MSE and MAE are defined as follows:
@@
$\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|z_{i}-\hat{z}_{i}|.$ (4)
Here, $n$ is the number of samples used for training, $z_{i}$ is the predicted value and $\hat{z}_{i}$ is the true value. The $R^{2}$ is a statistical measure that shows how well a model predicts the outcome in a regression analysis. It is the proportion between the variance explained by the model and the total variance:
@@
Here, $\bar{z}$ is the mean of the spectroscopic redshift values from the GAMA×$\times$×KiDS crossmatch. For an ideal case, the value of this metric would be unity.
The MSE is more sensitive to outliers than MAE. Due to the squaring part, MSE puts more weight on larger errors. MAE, on the other hand, treats all errors with equal importance, which can be advantageous when the dataset contains extreme values or noisy observations – which is common for galaxy images. Therefore, MSE is more appropriate for learning outliers, and MAE is better for ignoring them. To include the benefits of both these loss functions, we use the Huber (1964) loss function, similarly as in the previous KiDS analysis by Li et al. (2022). This function is quadratic when the absolute error is small, and linear when the absolute error exceeds a threshold $\delta$. This makes it robust to outliers because the impact of large errors is reduced compared to using a purely quadratic loss function such as MSE. The Huber loss is defined as
${L}(z,\hat{z})=\cases{}{1}{2}(z-\hat{z})^{2},&\text{if }|z-\hat{z}|\leq\delta,\\ \delta(|z-\hat{z}|-\frac{\delta}{2}),\text{otherwise}.{}$ (6)
@@
We evaluated the resulting photo-z performance using the following standard statistics:
@@
    $\delta z=z_{\mathrm{phot}}-z_{\mathrm{spec}};$ (7)
@@
    $\Delta z=\frac{\delta z}{1+z_{\mathrm{spec}}};$ (8)
  • standard deviation of normalized bias, $\sigma_{\Delta z}$;
@@ 4 Photometric redshift model
We used the Rectified linear unit (Relu, Jentzen et al., 2023) as the activation function in all the layers except in the output one where the sigmoid function (i.e., logistic curve) is used, which enforces all the predictions to lie in the range $10<z_{\mathrm{phot}}<$ 1. For the layers other than the last (output) one, we also tried other activation functions such as leaky Relu (Jentzen et al., 2023), peaky Relu, softmax, swish, and tanh, but Relu gave the best performance.
For the output layer, the sigmoid function gives the best results in our redshift range where practically all the galaxies have $z<1$ due to the $r<20$ mag flux limit. For instance, in the entire GAMA spectroscopic sample, objects with $z>1$ constitute less than 0.1% of the total and these are typically very bright and rare AGNs. For such sparse sources, empirical methods such as ours would not be able to render reliable redshift predictions unless some special approach to anomaly handling is taken. We therefore sacrificed a very small number of objects that lie at true $z>1$ to have $z_{\mathrm{phot}}<1$ in order to avoid a gross redshift overestimation for others, which could happen if the model had more freedom. The latter is for instance the case for ANNz2, where some of the photo-zs from B21 are predicted significantly above unity. Similar logic applies to photo-$z$s with non-physical predictions of $z<0$ which are equally avoided in our model, while were present in the ANNz2 results. We note, however, that both cases were rare already in B21. In the clean KiDS-Bright sample, there were 82 galaxies with $z_{ANNz2}>1$ and 444 with $z_{ANNz2}<0$.
Figure 2: Architecture of the Hybrid-z model, which employs four-band KiDS galaxy images together with nine-band KiDS+VIKING magnitudes for photo-z derivation. For the CNN part (left-hand side), the symbols used are as in Fig. 1, while inception modules are represented as hexagons. In the fully connected part (right-hand side), dense layers are given as blue rectangles, and the number of neurons is shown in parentheses.
@@
The features extracted by the convolutional and average pooling layers are passed through the Inception module before reaching the final dense layers. We found that the performance of the model improves with the inclusion of Inception, with notable improvement found in the reduction of the loss function value. We conducted tests with different numbers of Inception modules, ranging from three to five. Our findings suggest that using four Inception modules with varying numbers of filters is optimal for our data. The first inception module receives input from the Average Pooling 2D layer, while each subsequent inception module takes its input from the output of the preceding inception module. The initial inception modules extract low-level features and the final inception layers capture the global patterns in the image. Padding is applied before the convolutional and pooling operations for each Conv2D and AveragePooling2D layer using the ‘same’ padding (He et al., 2015). This ensures that the spatial dimensions (height and width) of the output feature maps are maintained as much as possible relative to the input. When the kernel is applied to the input, padding is added around the edges as necessary to keep the spatial dimensions consistent after each convolution and pooling operation.
Figure 3: Performance of the Huber loss function during the training and validation stage of the Hybrid-z.
As in Li et al. (2022) and Henghes et al. (2022), for our photo-z model, we combined two different types of networks, ”shallow” (fully connected) ANN, that we will denote as ordinary neural networks (ONNs) and CNN, hence the name Hybrid-z. There are $\sim$13.8 million trainable parameters in our model. An important aspect of the Hybrid-z model is the concatenation step. The nine-band magnitudes of galaxies are the additional information for the network. These are processed by dense layers. The flattened feature map and ONN output are combined via depth-wise concatenation. ONN output has 64 features. However, when we include images, the number of features increases to 12,064. Then this concatenated information is passed through final layers for the prediction of photo-zs. This significant increase in features for final dense layers after concatenation highlights the substantial contribution of the image data.
In Fig. 3, the performance of the Huber loss function for the Hybrid-z is shown. The addition of supplementary information enhances the model’s performance with respect to using only 4-band images. We find that for this model, we can achieve $R^{2}$ value greater than 0.93. We used early stopping criteria (Prechelt, 1996) for determining the number of epochs. The threshold for considering an improvement is zero, meaning that any decrease in the validation loss would count as an improvement. Early stopping will be triggered after ten consecutive epochs where the validation loss does not improve and the training will automatically stop.
@@ 5.1 Hybrid-z performance on test data
    General test data from the KiDS-DR4 Bright sample crossmatched with GAMA in the equatorial fields, including both clean and contaminated galaxies.
@@
Figure 4: Comparison of the true spectroscopic redshifts with the photometric ones derived with our framework, Hybrid-z, using four-band KiDS images and nine-band KiDS+VIKING magnitudes. The test data were taken from a subset of GAMA crossmatched with KiDS. The thick red line depicts the median while the thin lines show the scatter (SMAD) around the median.
Figure 5: Photometric redshift errors of Hybrid-z as a function of four quantities, from left to right, (a) photometric redshift as derived in this work; (b) true spectroscopic redshift from the clean sample; (c) apparent $r$-band magnitude; and (d) observed $u-g$ color. The thick solid red line represents the running median, while the thin red lines illustrate the scatter (SMAD), based on a blind test sample derived from GAMA.
The comparison of the main statistics of Hybrid-z versus ANNz2 results are provided in Table 1. Our new model employing jointly optical imaging and KiDS+VIKING magnitudes performs clearly better than ANNz2 which used 9-band magnitudes. In the first block of the Table, we provide statistics of a general test sample, which consists of GAMA galaxies crossmatched with KiDS-Bright including both ‘clean’ (masked=0) and ‘contaminated’ (masked=1) data, where the ‘masked’ flag was derived in B21 based on the KiDS-internal ‘MASK’ bit-wise column. The compared methods have comparable mean residuals – that is, they are similarly accurate – but Hybrid-z outperforms ANNz2 in its photo-$z$ precision, quantified as the scatter in $\Delta z$. Both in terms of standard deviation and SMAD, our combined CNN+ONN approach performs better, be it for the overall or clean (masked=0) test data. Hybrid-z provides about 20% smaller scatter than the ANNz2-based derivations of B21, where 9-band magnitudes were used. In the case of the fiducial clean data that are recommended for science, we achieve SMAD$(\Delta z)\simeq 0.014$, as compared to 0.018 in B21. Our result is at the same level as that by Treyer et al. (2024) where CNNs were used for photo-$z$s based on SDSS images at similar depths as in our case.
@@
In panels (c) and (d) of Fig. 5 we illustrate the behavior of photo-$z$ errors of our Hybrid-z model as a function of two observables: the $r$-band apparent magnitude and the observed $u-g$ color. The former is shown for the ‘auto’ flux measurement, which approximates the total light of a galaxy and was used by B21 to select the flux-limited KiDS-Bright sample. We observe slightly growing scatter (thin red lines indicating SMAD in Fig. 5 (c)) at the faintest end of the sample ($r\lesssim 20$). Otherwise, photo-$z$s are stable as a function of magnitude, with practically zero bias at most of the range. The $u-g$ color shown in panel (d) can be used to split galaxies into red and blue populations (similarly as $u-r$, Strateva et al. 2001). Indeed, we see the bimodality in the plot, with bluer galaxies to the left and redder to the right. As expected for photo-$z$s, the former has a wider scatter with respect to the true value than the latter. The difference is, however, not substantial, and again the median biases for the whole range of the $u-g$ color are very close to zero.
@@
We further quantify the performance of Hybrid-z for blue and red galaxies in the two bottom blocks of Table 1. Here, for a direct comparison between this work and B21, the selection of the two galaxy types was based on their intrinsic rather than observed properties, namely via their positioning on the absolute $r$-band magnitude versus rest-frame $u-g$ color diagram [^4] More specifically, we crossmatched our sample with galaxies labeled as red and blue in KiDS-Bright DR4. as discussed in B21. Similarly as for the full sample, also for these subsets, Hybrid-z provides considerable improvement in precision over ANNz2. Interestingly, this reduction in SMAD($\Delta z$) is greater for blue galaxies (by $\sim 22\%$) than for red ones ($\sim 17\%$). This is consistent with the fact that, as compared to the approaches relying solely on summary measurements such as magnitudes, the DL-based methods, which directly extract features from images to estimate redshifts, are expected to show larger improvement for galaxies with intricate morphologies such as spirals, that typically have bluer colors (e.g., Schuldt et al., 2021; Treyer et al., 2024). In contrast, for red galaxies – typically ellipticals with fewer features in images – the reduction in scatter for Hybrid-z is smaller than for the overall sample, although we would such as to emphasize that it still does perform better than ANNz2 for the same galaxy selection. In addition, it is worth noting that SMAD($\Delta z$) of 0.0154 for blue galaxies as obtained by Hybrid-z is even smaller than the same statistic of ANNz2 for red ones (cf. blocks #3 and #4 of Table 1).
@@
Figure 6: Comparison of photometric redshift error distributions between Hybrid-z (blue, this work) and ANNz2 (B21, red) shown as normalized counts based on the test sample results. Solid lines illustrate the best-fit Gaussian (with $\sigma=0.0145$ and $\mu=2.56\times 10^{-4}$, yellow) and generalized Lorentzian (Eq. 10, with $a=2.332$ and $s=0.0116$, black). The inset compares the histogram of Hybrid-z photo-z residuals with the fit models in logscale of the $y$-axis.
@@
The final comparison between ANNz2 and Hybrid-z performance is provided in Fig. 6, where we show the distributions of photo-$z$ errors $\Delta z$ as calibrated on the GAMA test data. As discussed in B21, following earlier work (e.g., Bilicki et al., 2014), such a distribution presents non-Gaussian features in the wings and is therefore better fit with a ”generalized Lorentzian” of the form [^5] It is interesting to note that also spectroscopic redshifts may present Lorentizan rather than Gaussian uncertainties, e.g., Yu et al. (2024).
@@
To summarize this subsection, the Hybrid-z model using CNNs on KiDS optical images together with ONNs on KiDS+VIKING magnitudes provides considerably better results than the previous 9-band ANNz2 derivations of B21. The improvement from ANNz2 to Hybrid-z is at the level of 20% reduction in photo-$z$ scatter (better precision) while maintaining almost zero bias (very good accuracy). Thanks to our new model, we observe smaller photo-$z$ errors for various subsamples of test data, and most remarkably for blue galaxies, where SMAD($\Delta z$) is lowered by about 22% from what was obtained in B21. As such, these are state-of-the-art results for the KiDS-Bright galaxies and in line with independent derivations for SDSS at the same depth (Treyer et al., 2024).